# Quality or Quantity? On Data Scale and Diversity in Adapting Large Language Models for Low-Resource Translation

### v1

@@ 4.2 Evaluation
To evaluate MT into the 11 American languages, we use AmericasNLP’23 validation sets (Ebrahimi et al., 2024) containing spa-X translation pairs. For the Indic pairs, we use the WMT 2023 test sets from the Indic track (Pal et al., 2023) which consist of eng-X pairs. We show test set statistics in Table 5. Given the absence of neural metrics for these languages, we evaluate using ChrF++ (Popović, 2017), since both the American and Indic languages are morphologically rich. We use SacreBLEU (Post, 2018) for computing this. We also report confidence intervals with bootstrap resampling (Koehn, 2004), which we implement for the multilingual setting by computing macro average across all languages for each resample, and then computing the mean and variance across all resamples. We report the standard deviation as the confidence interval.
@@ 4.3 Experimental Settings
For language-wise temperature sampling, we use $\tau=30$ for CPT and $\tau=80$ for SFT. We used a batch size of 8 and gradient accumulation every 16 steps. We used a learning rate of 1e-4, with a cosine scheduler and a warmup ratio of 3%. We train on bf16 precision. We use Llama-Factory (Zheng et al., 2024) for training and evaluating all models, with Deepspeed ZeRO3 (Rasley et al., 2020) for distributed training. For inference, we use a batch size of 16 with greedy decoding.
@@
LoRA only 4.46 18.80 18.32
LoRA + embeddings 9.58 16.12 16.28
@@
Table 7: chrF++ scores for spa-X LLM-MT in the American languages, for LLMs of varying sizes. Both LoRA modules and embeddings are fine-tuned. For SFT, all models use 500K spa-X MT examples. Confidence estimates are computed using bootstrap resampling.
@@ 4.4 Foundational Results
We first establish the importance of fine-tuning input and output embeddings in Table 6, and then, show our foundational results for the American languages in Table 7, using 5-shot prompting (Brown et al., 2020) as a baseline. Our findings are:
@@
    SFT alone is effective, but CPT+SFT yields best results. While SFT yields large gains over prompting, combining CPT and SFT seems optimal for both Gemma and Llama3, although the gap is smaller than that of high-resource MT (Xu et al., 2024), ostensibly due to the difference in scale. For Mistral, SFT alone seems to suffice - we hypothesize that this might be due to the smaller vocabulary being effectively fine-tuned on SFT data alone.
@@
    pre-training on parallel data yields major gains consistently. Lastly, we observe that pre-training on a mixture of concatenated monolingual and parallel data (‘mono + parallel’) yields statistically significant gains over using monolingual data alone (ie. ‘all mono’; refer Section 3.1). This trend is consistent for all 3 LLMs, with larger ones for the more effective models Mistral and Llama3. We note that for ‘mono+parallel’, we mix parallel and monolingual data in a 1:1 ratio, since we observed it worked best empirically, and show in Table 14 how increasing the ratio of parallel data during CPT monotonically improves performance. Given these exciting gains, we explore the importance of parallel data for low-resource LLM-MT further in Section 4.5.
@@ 4.5 Analysis: Importance of Parallel Data
Having seen improvements in pre-training on the entire corpus (consisting of 13M ‘mono+parallel’ sentences, or 730M tokens) in Table 7, we now study the importance of parallel data as we scale down – an important consideration when one moves to even lower-resource settings. We pretrain on subsets of varying sizes, and mix monolingual and parallel data in 3 ways: ‘All mono’, ‘mono+parallel (concat)’ and ‘mono+parallel (separate)’, as defined in Section 3.1). We fine-tune all these pre-trained models on the same SFT dataset: 500K spa-X MT instructions, and plot the resulting chrF++ scores in Figure 2, including error bars from bootstrap resampling. We note that ‘all mono’ has different markers than the others, as the token counts on including both source and target-side data in the corpus are obviously larger than only the latter. We find that, a) starting at around 5M sentences (~300M tokens), we observe that it is consistently advantageous to include concatenated parallel data during pre-training. b) Given ‘mono+parallel (separate)’ severely underperforms, we establish that it is concatenation that adapts the LLM for the task of MT, not the extra data alone. Our findings complement those of Alves et al. (2024), who also observe gains from pre-training on parallel data in the 1B to 20B tokens range, using 10 high-resourced European languages. In contrast, our focus here is on investigating the minimum data threshold at which CPT leveraging parallel data becomes beneficial.
Figure 2: Comparing Llama3 8B models pre-trained on monolingual data alone versus those that include parallel data too - concatenated, or as separate texts - at various scales. All models were pre-trained on 1M, 3M, 5M, 8M, and 13M sentences respectively, and markers denote the corresponding token counts. The Y-axis shows chrF++ post SFT on 500K spa-X MT data for 1 epoch.
@@
Figure 3: Scaling up Llama3 8B models with different CPT recipes (no CPT, CPT with monolingual data, and CPT with a mixture of monolingual and parallel data) on MT data for the American languages. ‘High-resource’ refers to the relatively higher-resourced languages in our low-resource setup (Aymara, Guarani and Quechua) while the other 8 are grouped as low-resourceη
Figure 4: Scaling up Llama 3 8B models with the 3 CPT recipes for the 4 medium-resourced Indic languages, until 5M sentences. We were forced to stop training ‘No CPT’ at 2.5M sentences due to budget constraints.
@@
We now turn our focus to scaling during SFT, which has not yielded gains in high-resource LLM-MT after 20K-30K sentences (Zhang et al., 2024a). Motivated by our previous results, we re-examine this question for low-resource LLM-MT in Figure 3, wherein we evaluate at steps of 50K until ~1M SFT instructions - when our spa-X MT dataset runs out. Our findings are as follows:
@@
    Overall MT quality improves steadily with scale. Unlike high-resource LLM-MT, we observe that, despite fluctuations, MT quality continues to grow until 1M sentences – particularly for LRLs, which have a steeper slope.
@@
    LRLs need SFT scaling much more than HRLs So, while HRLs benefit hugely from CPT on bitext, the opposite is true for LRLs – the scarce amount of parallel data observed during CPT (see Table 3) is likely not enough to outperform ‘all mono’ pre-training. Instead, LLMs are far more responsive to the scale of LRL MT data during SFT, showing consistent performance improvement. Meanwhile, MT quality plateaus for the HRLs333although Aymara, Guarani and Quechua are LRLs for the original pre-training corpus of Mistral/Llama3 (Figure 3(b)) - suggesting that the ‘less is more’ (Zhou et al., 2023) trend popular in high-resource LLM-MT does not hold for LRLs, where scale continues to remain the most effective option.
For generalizability, we also provide scaling graphs for Mistral 7B in Figure 7 and report trends similar to Llama3. Moreover, we also compute scaling graphs for the 4 medium-resourced Indic languages in Figure 4 using the Llama3 model, until 5M eng-X sentences. Here, we observe that the gap between ‘CPT: mono+parallel (concat)’ and ‘CPT (all mono)’ is relatively lesser, and only significant until 500K sentences, at which point they start becoming comparable. This might be because the Indic parallel corpora used for CPT is relatively less diverse, consisting of only eng-X pairs in 4 languages, whereas our parallel data for the American languages is more heterogeneous (see Table 3) – which we observed to be yield more gains in our preliminary experiments. Thus, we expect a more diverse corpus for the Indic languages to yield greater long-term gains, but we leave the verification of this to future work. In other respects, the trends are quite similar to the American languages – the gains of CPT carry over to SFT here too and, interestingly, MT quality continues to improve until 5M SFT examples - once again in direct contrast to high-resourced LLM-MT research.
@@
Finally, given the consistent gains from scale, we evaluate how effective data repetition could be. We train for up to 10 epochs on the entire 1M spa-X MT dataset and plot our results in Figure 5. We observe a strong monotonic gain of +3.3 average chrF++ until 5 epochs, with the largest coming from the 1st to the 2nd epoch (+2.0 chrF++). While the graph does plateau later (likely due to overfitting) simply training for more epochs is an easy way to boost performance, given the data constraints for LRLs.
@@ 4.6 Analysis: Diversity in SFT is not always helpful
We now look at the importance of diversity in SFT, which plays a significant role in boosting LLM performance during general-purpose fine-tuning (Sanh et al., 2022; Longpre et al., 2023), but has also been favoured in training LLM-MT systems (Alves et al., 2024). Alves et al. (2024) observed improved performance on non-MT tasks on including non-MT instructions, but mixed results on MT quality - concluding that the “transfer/interference relations between tasks are complex”. In this work, we try to tease apart these interference relations, and study the effect of non-MT tasks on low-resource translation in depth. In particular, we look at 3 kinds of diversity: prompt, linguistic, and task diversity, and examine their impact on MT quality when mixed in varying proportions in the SFT dataset.
@@
Table 9: Exploring interference due to diversity in SFT for our best Llama3 8B model (CPT on both monolingual and parallel data). The dataset size (example count) is prepended before each task. Scores shown are average chrF++ across all 11 languages.
@@
Inspired by the previous results showing target-side diversity in MT pairs boosts performance, we now broaden the scope of diversity during SFT. In Table 9, we show varying mixtures of SFT data that all have the same size (500K examples) but are composed of different tasks. For prompt diversity, we ablate randomly sampling from a list of potential prompt templates (listed in Table 11) versus using a constant one (the first in Table 11). We observe a statistically significant gain in MT quality, similar to general-purpose SFT (Longpre et al., 2023).
Next, we study the more interesting question of linguistic diversity – can data in other MT pairs transfer for test languages? In one baseline, we divide our SFT dataset equally into spa-X, eng-X, and por-X MT examples, into the American languages (statistics in Table 3). In another baseline, we use Spanish-English data from ParaCrawl (Bañón et al., 2020) and combine a 250K sentence subset with 250K pairs from our usual spa-X MT data. We find that this type of linguistic diversity leads to interference, and significantly underperforms the 500K spa-X MT baseline. This suggests that while target diversity in related languages might help performance, source diversity or unrelated high-resource languages like English may not.
Finally, we look at task diversity - can non-MT tasks that elicit better instruction-following and general reasoning capabilities in the source or target language, benefit LLM-MT in LRLs? We mix Aya and XQA (Section 3.2) instructions with a 250K subset of spa-X MT examples. We also include an ablation that, in place of curating non-MT data, merely trains for 2 epochs on the 250K spa-X MT subset. We discover that, here too, general-purpose tasks lead to interference, and training for 2 epochs on 250K MT examples is a more effective strategy that is comparable to 1 epoch on 500K.
Figure 6: Mixing varying percentages of SFT tasks with spa-X MT, versus impact on spa-X chrF++, for 500K total examples. ‘MT only’ is a baseline which uses only spa-X MT data from the other plots, discarding all else. Thus, it uses less data, and helps to estimate interference.
@@
Finally, we seek to establish conclusively if the negative results of diversity in Table 9 arise from factors like task mismatch or lack of quality; or if interference is simply the norm when fine-tuning on non-MT data for low-resourced LLMs. In Figure 6, we plot 5 graphs: the first 4 are graphs showing non-task-specific data (including XQA, Aya, {eng/por}-X MT and eng-spa MT) mixed in varying proportions with spa-X MT. Thus, for example, the second point of each plot represents 10% of that task mixed with 90% of spa-X MT, with the total always being 500K examples. We also include an ablation (shown in dotted lines), which has the same quantity of spa-X MT data at each data point as the other plots, only without data from any of the other tasks listed. Thus, the second point of this ablation contains 450K sentences, the third one 375K, and so on.
We observe that performance is always better using task-specific MT data, and negative interference is indeed consistent. We conclude that transfer is challenging to achieve in low-resource settings, likely because at such scales, the LLM-MT model is still learning to generate translations, and reasoning is still a formidable challenge. Downstream performance in such settings depends not on diversity/composition of the SFT dataset, but only on the amount of LRL MT data provided – making optimizing for quantity the most effective strategy.

### v2

@@ 4.2 Evaluation
To evaluate MT into the 11 American languages, we use AmericasNLP’23 validation sets (Ebrahimi et al., 2024) containing spa-X translation pairs. For the Indic pairs, we use the WMT 2023 test sets from the Indic track (Pal et al., 2023) which consist of eng-X pairs. Both evaluation datasets are multi-domain, as are the curated monolingual and parallel corpora. We show test set statistics in Table 5. Given the absence of neural metrics for these languages, we evaluate using ChrF++ (Popović, 2017), since both the American and Indic languages are morphologically rich wherein chrF++ is particularly effective (Popović, 2017). We use SacreBLEU (Post, 2018) for computing this. We also report confidence intervals with bootstrap resampling (Koehn, 2004), which we implement for the multilingual setting by computing the macro-average across all languages for each resample, and then computing the mean and variance across all resamples. We report the standard deviation as the confidence interval.
@@ 4.3 Experimental Settings
For temperature sampling our data, we use $\tau=30$ for CPT and $\tau=80$ for SFT. We used a batch size of 8 and gradient accumulation every 16 steps. We used a learning rate of 1e-4, with a cosine scheduler and a warmup ratio of 3%. We train all models on bf16 precision for 1 epoch. We use Llama-Factory (Zheng et al., 2024) for training and evaluating all models, with Deepspeed ZeRO3 (Rasley et al., 2020) for distributed training. For inference, we used a batch size of 16 with greedy decoding, since we found higher beam sizes yielded minimal gains.
@@
LoRA only 4.5 ± 0.1 18.8 ± 0.2 18.3 ± 0.2
LoRA + embeddings 9.6 ± 0.3 15.6 ± 0.4 15.6 ± 0.4
@@
Table 7: chrF++ scores for spa-X LLM-MT in the American languages, for LLMs of varying sizes. Both LoRA modules and embeddings are fine-tuned. For SFT, all models use 500K spa-X MT examples. Confidence estimates are computed using bootstrap resampling. ‘mono+parallel’ uses concatenated bitext.
@@ 4.4 Foundational Results
We first establish the importance of fine-tuning input and output embeddings in Table 6 and then show our foundational results for the American languages in Table 7, using 5-shot prompting as a baseline. Our findings are:
@@
    SFT alone is effective, but CPT+SFT yields best results. While SFT yields large gains over prompting, combining CPT and SFT seems optimal for both Gemma and Llama3, although the gap is smaller than that of high-resource MT (Xu et al., 2024), ostensibly due to the difference in scale. For Mistral, SFT alone seems to suffice—we hypothesize that this might be due to the smaller vocabulary being effectively fine-tuned on SFT data alone.
@@
    Pre-training on parallel data yields major gains consistently. Lastly, we observe that pre-training on a mixture of concatenated monolingual and parallel data (‘mono + parallel’) yields statistically significant gains over converting both as monolingual data (ie. ‘all mono’; refer Section 3.1). This trend is consistent for all 3 LLMs, with larger gains for the more effective models Mistral and Llama3. We note that for ‘mono+parallel’, we mix parallel and monolingual data in a 1:1 ratio, since we observed it worked best empirically, and show in Table 14 how increasing the ratio of parallel data during CPT monotonically improves performance. Given these gains, we explore the importance of bitext for low-resource LLM-MT further in Section 4.5.
@@ 4.5 Analysis: Importance of Parallel Data
Having seen improvements in pre-training on the entire corpus (consisting of 13M ‘mono+parallel’ sentences, or 730M tokens) in Table 7, we now study the importance of parallel data as we scale down—an important consideration when one moves to even lower-resource settings. We pre-train on subsets of varying sizes and mix monolingual and parallel data in 3 ways: ‘All mono’, ‘mono+parallel (concat)’ and ‘mono+parallel (separate)’, as defined in Section 3.1). We fine-tune all these pre-trained models on the same SFT dataset: 500K spa-X MT instructions, and plot the resulting chrF++ scores in Figure 2, including error bars from bootstrap resampling. We note that ‘all mono’ has different markers than the others, as the token counts on including both source and target-side data in the corpus are obviously larger than only the latter. We find that a) starting at around 5M sentences (~300M tokens), it is consistently advantageous to include concatenated parallel data during pre-training. b) Given ‘mono+parallel (separate)’ severely underperforms, we establish that it is concatenation that adapts the LLM for the task of MT, not the extra data alone. Our findings complement those of Alves et al. (2024), who also observe gains from pre-training on parallel data in the 1B to 20B tokens range, using 10 high-resourced European languages. In contrast, our focus here is on investigating the minimum data threshold at which CPT leveraging parallel data becomes beneficial.
Figure 2: Comparing Llama3 8B models pre-trained on monolingual data alone versus those included parallel data too—concatenated, or as separate texts at various scales. All models were pre-trained on 1M, 3M, 5M, 8M, and 13M sentences respectively, and markers denote the corresponding token counts. The y-axis shows chrF++ post SFT on 500K spa-X MT data for 1 epoch.
@@
Figure 3: Scaling up Llama3 8B models with different CPT recipes (no CPT, CPT with monolingual data, and CPT with a mixture of monolingual and parallel data) on MT data for the American languages. ‘High-resource’ refers to the relatively higher-resourced languages in our low-resource setup (Aymara, Guarani and Quechua) while the other 8 are grouped as low-resourceη.
Figure 4: Scaling up Llama 3 8B models with the 3 CPT recipes for the 4 Indic languages, until 5M sentences. We were forced to stop training ‘No CPT’ at 2.5M sentences, constrained by budget.
@@
We now turn our focus to scaling during SFT, which has not yielded gains in high-resource LLM-MT after 20K-30K sentences (Zhang et al., 2024a). Motivated by our previous results, we re-examine this question for low-resource LLM-MT in Figure 3, wherein we evaluate at steps of 50K until ~1M SFT instructions, the point at which spa-X MT data runs out. Our findings are:
@@
    Overall MT quality improves steadily with scale. Unlike high-resource LLM-MT, we observe that, despite fluctuations, MT quality continues to grow until 1M sentences, particularly for LRLs, which have a steeper slope.
@@
    LRLs need SFT scaling much more than HRLs So, while HRLs benefit hugely from CPT on bitext, the opposite is true for LRLs—the scarce amount of parallel data observed during CPT (see Table 3) is likely not enough to outperform ‘all mono’ pre-training. Instead, LLMs are far more responsive to the scale of LRL MT data during SFT, showing consistent performance improvement. Meanwhile, MT quality plateaus for the HRLs333although Aymara, Guarani and Quechua are LRLs for the original pre-training corpus of Mistral/Llama3 (Figure 3(b)), suggesting that the ‘less is more’ (Zhou et al., 2023) trend popular in high-resource LLM-MT does not hold for LRLs, where scale continues to remain the most effective option.
For generalizability, we also provide scaling graphs for Mistral 7B in Figure 7 and report trends similar to Llama3. Moreover, we also compute scaling graphs for the 4 Indic languages in Figure 4 using the Llama3 model, until 5M eng-X sentences. Here, we observe that the gap between ‘CPT: mono+parallel (concat)’ and ‘CPT (all mono)’ is relatively lesser, and only significant until 500K sentences, at which point they start becoming comparable. This might be because the Indic parallel corpora used for CPT is relatively less diverse, consisting of only eng-X pairs in 4 languages, whereas our parallel data for the American languages is more heterogeneous (see Table 3)—which we observed to yield more gains in our preliminary experiments. Thus, we expect a more diverse corpus for the Indic languages to yield greater long-term gains, but we leave the verification of this to future work. In other respects, the trends are quite similar to the American languages: the gains of CPT carry over to SFT here too and interestingly, MT quality continues to improve until 5M SFT examples, once again in direct contrast to high-resourced LLM-MT research.
@@
Finally, given the consistent gains from scale, we evaluate how effective data repetition could be during SFT. We train for up to 10 epochs on the entire 1M spa-X MT dataset and plot our results in Figure 5. We note a strong monotonic gain of +3.3 average chrF++ until 5 epochs, with the largest coming from the 1st to the 2nd epoch (+2.0 chrF++). While the graph does plateau later (likely due to overfitting) this shows how simply training for more epochs can be an easy way to boost performance, given the data constraints for LRLs.
@@ 4.6 Analysis: Diversity in SFT is not always helpful
We now look at the importance of diversity in SFT, which plays a significant role in boosting LLM performance during general-purpose fine-tuning (Sanh et al., 2022; Longpre et al., 2023), but has also been favoured in training LLM-MT systems (Alves et al., 2024). Alves et al. (2024) observed improved performance on non-MT tasks by including non-MT instructions, but mixed results on MT quality—concluding that the “transfer/interference relations between tasks are complex”. In this work, we try to tease apart these interference relations and study the effect of non-MT tasks on low-resource translation in depth. In particular, we look at 3 kinds of diversity: prompt, linguistic, and task diversity, and examine their impact on MT quality when mixed in varying proportions in the SFT dataset.
@@
(a) Prompt, Linguistic and Task Diversity in American pairs
SFT Mixture chrF++
250K spa-X MT + 250K Aya (asm, mni) 26.91 ± 0.42
250K spa-X MT + 250K Alpaca (eng) 28.02 ± 0.41
500K spa-X MT only 29.57 ± 0.43
(b) Task Diversity in Indic pairs
Table 9: Exploring interference due to diversity in SFT for our best Llama3 8B model (CPT on both monolingual and parallel data) on the American and Indic languages. The dataset size (example count) is prepended before each task. Scores shown are average chrF++.
@@
Inspired by the previous results showing target-side diversity in MT pairs boosts performance, we now broaden the scope of diversity during SFT. In Table 9(a), we show varying mixtures of SFT data that all have the same size (500K examples) but are composed of different tasks. For prompt diversity, we ablate randomly sampling from a list of potential prompt templates (listed in Table 11) versus using a constant one (the first in Table 11). We observe a statistically significant gain in MT quality, similar to general-purpose SFT (Longpre et al., 2023).
Next, we study the more interesting question of linguistic diversity: can data in other MT pairs transfer for test languages? In one baseline, we divide our SFT dataset equally into spa-X, eng-X, and por-X MT examples, into the American languages (statistics in Table 3). In another baseline, we use Spanish-English data from ParaCrawl (Bañón et al., 2020) and combine a 250K sentence subset with 250K pairs from our usual spa-X MT data. We find that this type of linguistic diversity leads to interference, and significantly underperforms the 500K spa-X MT baseline. This suggests that while target diversity in related languages might help performance, source diversity or unrelated high-resource languages like English may not.
Then, we look at task diversity: can non-MT tasks that elicit better instruction-following and general reasoning capabilities in the source or target language, benefit LLM-MT in LRLs? We mix Aya and XQA (Section 3.2) instructions with a 250K subset of spa-X MT examples. We also include an ablation that, in place of curating non-MT data, merely trains for 2 epochs on the 250K spa-X MT subset. Here we also discover that general-purpose tasks lead to interference and that training for 2 epochs on 250K MT examples is a more effective strategy comparable to 1 epoch on 500K.
We also look at the impact of task diversity on the Indic languages in Table 9(b). Here, we are able to find general-purpose instruction-tuning data in Assamese and Meitei, as mentioned in Section 3.2. Interestingly, including data in these target languages—the languages we want the model to be better at generating in—degrades performance the most! A similar amount of Alpaca data in English reduces performance far less, suggesting that reasoning in LRLs is such a hard task, that even data where the model should learn to generate in the languages of interest, is worse than generic English instruction-following data. Here, too, providing 500K spa-X examples is the most optimal strategy, underscoring the generalizability of our findings.
Figure 6: Mixing varying percentages of SFT tasks with spa-X MT, versus the impact on spa-X chrF++, for 500K examples. ‘MT only’ is a topline which only uses the spa-X MT data excluding all other tasks. Thus, it uses lesser data relatively and helps to estimate the interference of non-MT tasks on MT quality.
@@
Finally, we seek to establish conclusively if the negative results of diversity in Table 9(a) arise from factors like task mismatch or lack of quality; or if interference is simply the norm when fine-tuning on non-MT data for low-resourced LLMs. In Figure 6, we plot 5 graphs: the first 4 are graphs showing non-task-specific data (including XQA, Aya, {eng/por}-X MT and eng-spa MT) mixed in varying proportions with spa-X MT. Thus, for example, the second point of each plot represents 10% of that task mixed with 90% of spa-X MT, with the total always being 500K examples. We also include an ablation (dotted lines), which has the same quantity of spa-X MT data at each data point as other plots but without data from any other task listed.
We observe that performance is always better using task-specific MT data, and negative interference is indeed consistent. We conclude that transfer is challenging to achieve in low-resource settings, likely because at such scales, the LLM-MT model is still learning to generate translations, and reasoning is still a formidable challenge. Downstream performance in such settings depends not on the diversity/composition of the SFT dataset, but only on the amount of LRL MT data provided—making optimizing for quantity the most effective strategy.